A proton is moving with velocity 6.7 x 10^4 m/s i + 3.6 x 10^4 m/s j in a magnetic field produced by the current-carrying wire at a point 5.6 cm from the wire. The current 6.7 A is passing through the wire in an upward direction. Find the acceleration of the proton
1
Expert's answer
2019-10-07T10:57:37-0400

When moving under the action of the Lorentz force

"a = \frac {2\cdot 10^{-7}qv_nI}{mr} = 10^7\frac {2\cdot1.6\cdot 6.7\cdot 6.7}{1.67\cdot 5.6} = 154(Mm/s^2)"